### v1

@@ 3.1 Some properties of the solution operator
  • For each $i$ and $j$, $\mathcal{P}_{i,j}$ is an operator whose range lies within the generalized eigenspace corresponding to the eigenvalue $\lambda_{i}$; for each $i$, $\mathcal{P}_{i,m_{i}}$ is nonzero.
@@
Based on these two properties, a simple analytic-continuation argument shows that, given a real interval $[k_{\textup{min}},k_{\textup{max}}]\subset\mathbb{R}\setminus\{0\}$, there exists a finite-thickness “strip”
@@ 3.2 Field amplification
Computing $\phi$ in practical settings demands a discretization of the infinite-dimensional space $[L^{2}(\Gamma)]^{2}$. As such, we replace the analytic field amplification Eq. 14 with its discrete counterpart, which we still denote by $\phi$ with an abuse of notation:
@@
Let $\lambda\in\mathbb{C}$ be such that $\textup{A}(\lambda)$ is singular, i.e., $\lambda\in\Lambda$. Let $m=m(\lambda)\in\mathbb{N}$ be the length of the longest chain of generalized eigenvectors associated to $\lambda$, cf. Remark 3.2. As $k\to\lambda$, there exist $\psi_{\lambda}>0$, depending only on the spectral projector $\textup{P}_{1,m}$, cf. Theorem 3.1, such that
@@
This case is typical in practical settings, e.g., since round-off noise makes all poles’ orders 1.
@@ 4.1 Rational approximation
In order to justify rational approximation for field amplification, we take Theorem 3.1 as starting point. Therein, it is shown that the resolvent $\textup{A}(k)^{-1}$ is a matrix-valued meromorphic function of the wavenumber $k$. As such, it is possible to efficiently approximate it to arbitrary accuracy via matrix-valued rational functions, which, in this context, we define as the ratio of a matrix-valued polynomial and a scalar polynomial. For instance, using the extremely popular barycentric interpolatory rational form [24], we may approximate $\textup{C}(k)$ as
@@
Note how the rational approximation consists of a linear combination of samples $\{\textup{C}(\overline{k}_{i})\}_{i=1}^{n_{K}}$ at arbitrary (distinct) locations $\{\overline{k}_{i}\}_{i=1}^{n_{K}}\subset\mathbb{C}$, weighted by scalar-valued rational functions
@@ 4.2 Compression by randomized power method
This is simply the square root of the estimate for the largest eigenvalue of $^{*}\textup{C}$ C that one gets after $q$ iterations of the well-known power method with starting vector $\mathbf{b}$.
@@ 5 Hybrid surrogate modeling
In the previous sections, we have discussed two surrogate-modeling-based methods for approximating $\phi$, which follow rigorous function-approximation and computational-linear-algebra theory. In this section, we describe an alternative approach, which approximates $\phi$ following more heuristic arguments. Thanks to some drastic simplifications, the upcoming “hybrid” method, described in Section 5.3 below, is extremely efficient. Indeed, it builds a surrogate model using roughly as few samples as Algorithm 2, with an extremely low surrogate-evaluation cost $c_{\textup{surrogate}}=\mathcal{O}\left(n_{K}\right)$.
@@ 5.2 Assembling the approximate field amplification
We start by observing that the coefficients $\{\psi_{\lambda}\}_{\lambda\in\widetilde{\Lambda}_{K}}$ can be interpreted as the degrees of freedom appearing in an approximation of $\phi$ that uses a linear combination of the basis functions $\{\left|\cdot-\lambda\right|^{-1}\}_{\lambda\in\widetilde{\Lambda}_{K}}$. As long as the real parts of the approximate poles are pairwise distinct, any set of $M_{K}^{\prime}$ distinct collocation points $\{\widetilde{k}_{i}\}_{i=1}^{M_{K}^{\prime}}\subset\mathbb{R}$ is “unisolvent” with this basis, in the sense that the data-fit problem
@@
Leveraging common ideas from the radial-basis literature, one may improve both approximation accuracy and numerical stability by placing the collocation points at the locations where the basis functions attain their respective maxima. See, e.g., [34]. In our setting, this happens at $\widetilde{k}_{i}:=\operatorname*{Re}\lambda_{i}$ for $i=1,\ldots,M_{K}^{\prime}$.
@@
The numerators $\widetilde{\psi}_{\lambda}$ may be found, e.g., with a heuristic collocation-based approach, through a direct sampling of $\phi(\widetilde{k})$ at $\widetilde{k}=\operatorname*{Re}\lambda$:
@@
By construction, this simple choice yields $\phi(\operatorname*{Re}\lambda)=\widetilde{\phi}_{\max,0}(\operatorname*{Re}\lambda)$ for any $\lambda\in\Lambda_{K}$, as long as the maximum in Eq. 30 is attained at $\lambda$, i.e., whenever no approximate pole “hides” any other approximate pole.
@@ 5.3 The overall algorithm
     Evaluate $\phi(\widetilde{k})${cost $\mathcal{O}\left(N_{h}^{3}\right)$}
@@ 6.1.1 Spectral features
Moreover, we can see that one of the rational poles, $\lambda\approx 3.16+0.07\textup{i}$, has positive imaginary part. This is a non-physical eigenvalue, but should not be necessarily considered spurious, cf. Remark 5.1. Indeed, rational approximation sometimes relies on unstable poles to improve approximation accuracy. It is commonly suggested to either keep such non-physical poles, to “flip” them [13], i.e., change the sign of their imaginary parts, or to remove them [18]. Four more spurious poles can also be observed in Fig. 2 (top), along a band with imaginary part between $-1$ and $-0.2$.
@@ 6.1.2 Field amplification
A partial explanation for this effect can be found by looking at the shape of the $14$ basis functions $\left|\cdot-\lambda\right|^{-1}$, cf. Section 5.2, that are removed when poles are filtered. As a specific example, the basis function corresponding to the spurious pole $\lambda_{1}\approx 1.20-0.62\textup{i}$, namely, $((k-1.20)^{2}+0.62^{2})^{-1/2}$, is more “spread out” than the basis function corresponding to the non-spurious pole $\lambda_{2}\approx 1.22-0.04\textup{i}$, whose real part is similar but whose imaginary part is much smaller, namely, $((k-1.22)^{2}+0.04^{2})^{-1/2}$. Due to its larger variance, the former basis function has a global effect on the approximation $\widetilde{\phi}_{\textup{sum},0}$, breaking the “upper-bound” property at wavenumbers far from $\operatorname*{Re}\lambda_{2}$.
We also test the hybrid algorithm with $n_{b}>0$ extra basis elements, by taking a piecewise-linear finite element basis on a uniform partition of $K$ into $(n_{b}-1)$ intervals. In our experiment, we set $n_{o}=n_{b}=M_{K}^{\prime}=20$, the number of (filtered) rational poles. This leads to an enhanced approximation accuracy. Interestingly, the “max” approximation $\widetilde{\phi}_{\textup{max},n_{b}}$, cf. Eq. 30, seems to perform better than the “sum” one $\widetilde{\phi}_{\textup{sum},n_{b}}$, cf. Eq. 27.
@@ 6.2 C-shaped and kite-shaped scatterer
C shape Algorithm 1 $3.60\cdot 10^{4}$
@@
kite Algorithm 1 $3.21\cdot 10^{4}$
@@ 7 Conclusions
  • The “hybrid” method from Algorithm 4 targets the field amplification more directly, through a scalar interpolatory approach. This reduces the evaluation cost at the price of some accuracy. The hybrid method uses rational approximation as a tool to greatly improve the approximation quality and efficiency, by enriching the interpolation basis with some additional rational terms. This strategy ultimately enables a modest approximation accuracy at extremely low cost.
@@
Concerning future research directions, we note that the field amplification has been defined in terms of surface potentials on the interface $\Gamma=\partial\Omega$. In some cases, it might be of greater interest to consider volume field amplifications, e.g., measured in the $H^{1}(\Omega)$-metric: $\phi_{\Omega}(k):=\sup_{u_{\textup{inc}}\neq 0}\left\|u_{\textup{tot}}(k)\right\|_{H^{1}(\Omega)}/\left\|u_{\textup{inc}}\right\|_{H^{1}(\Omega)}$, cf. Eq. 1. In our setting, $\phi_{\Omega}$ may be computed via the representation formula Eq. 5, which provides an invertible extension operator from surface potentials to Helmholtz solutions over $\mathbb{R}^{d}$. In principle, our discussion and our techniques should apply also to $\phi_{\Omega}$. However, in such case the approximation task is much more challenging, since (i) the wavenumber-dependent extension operator makes $\phi_{\Omega}$ depend on $k$ in a more complicated way than $\phi$, and (ii) building and inverting the extension operator is difficult in itself. For these reason, further analysis and implementation work remain to be done.
@@ Appendix A Exact scattering eigenvalues on the disk
find $\textup{find }\lambda\in\mathbb{C}\textup{ and }(\alpha,\beta)\in\mathbb{C}^{2}\setminus\{\mathbf{0}\}\textup{ s.t.}\begin{pmatrix}J_{\nu}(n_{\textup{i}}\lambda)&H_{\nu}(\lambda)\\ n_{\textup{i}}\lambda J_{\nu}^{\prime}(n_{\textup{i}}\lambda)&\lambda H_{\nu}^{\prime}(\lambda)\end{pmatrix}\begin{bmatrix}\alpha\\ -\beta\end{bmatrix}=\mathbf{0}.$

### v2

@@ 3.1 Some properties of the solution operator
  • For each $i$ and $j$, $\mathcal{P}_{i,j}$ is an (oblique) projection operator whose range lies within the generalized eigenspace corresponding to the eigenvalue $\lambda_{i}$; for each $i$, $\mathcal{P}_{i,m_{i}}$ is nonzero.
@@
Based on these two properties, a simple analytic-continuation argument shows that, given a real interval not containing 0, $k_{\textup{min}},k_{\textup{max}}]\subset\mathbb{R}\setminus\{0\}$, there exists a finite-thickness “strip”
@@ 3.2 Field amplification
Computing $\phi$ in practical settings demands a discretization of the infinite-dimensional space $L^{2}(\Gamma)]^{2}$. As such, we replace the analytic field amplification Eq. 14 with its discrete counterpart, which we still denote by $\phi$ with an abuse of notation:
@@
Let $\lambda\in\mathbb{C}$ be such that $\textup{A}(\lambda)$ is singular, i.e., $\lambda\in\Lambda$. Let $m=m(\lambda)\in\mathbb{N}$ be the length of the longest chain of generalized eigenvectors associated to $\lambda$, cf. Remark 3.2. As $k\to\lambda$, there exist $\psi_{\lambda}>0$, depending only on the mass matrix M and on the spectral projector $\textup{P}_{1,m}$, cf. Theorem 3.1, such that
@@
The above assumption is typical in practical settings, since, even if defective poles ($m(\lambda)>1$) exist, general perturbations of the scatterer $\Omega$ make the poles (semi-)simple again, recovering $m(\lambda)=1$ [36]. From a “numerical” perspective, this means that round-off noise almost surely makes all poles’ orders 1.
More generally speaking, in cases where the spectrum $\Lambda$ includes defective eigenvalues ($m(\lambda)>1$), one may want to modify our upcoming techniques to account for the proper exponent m-m in Eq. 16. Still, in all our numerical experiments, we have found no evidence to discard the assumption that $m(\lambda)=1$ for all $\lambda$. This fact, together with the intrinsic difficulties related to numerically dealing with defective eigenvalues (notably, identifying $m(\lambda)$), behooves us to leave the “defective case” as future research.
@@ 4.1 Rational approximation
In order to justify rational approximation for field amplification, we take Theorem 3.1 as starting point. Therein, it is shown that the resolvent $\textup{A}(k)^{-1}$ is a matrix-valued meromorphic function of the wavenumber $k$. As such, it is possible to efficiently approximate it to arbitrary accuracy via matrix-valued rational functions, which, in this context, we define as the ratio of a matrix-valued polynomial and a scalar polynomial. For instance, using the barycentric interpolatory rational form [24], we may approximate $\textup{C}(k)$ as
@@
Barycentric rational interpolation has been extremely popular in recent years, thanks to its beneficial stability properties even for scattered (non-structured) sets of interpolation points $\{\overline{k}_{i}\}_{i=1}^{n_{K}}\subset\mathbb{C}$, which makes it superior to many comparable approximation techniques. Note how the rational approximation consists of a linear combination of samples $\{\textup{C}(\overline{k}_{i})\}_{i=1}^{n_{K}}$ at arbitrary (distinct) locations $\{\overline{k}_{i}\}_{i=1}^{n_{K}}$, weighted by scalar-valued rational functions
@@ 4.2 Compression by randomized power method
This is simply the square root of the estimate for the largest eigenvalue of $\textup{C}^{*}\textup{C}$ that one gets after $q$ iterations of the well-known power method with starting vector $\mathbf{b}$.
@@ 5 Hybrid surrogate modeling
In the previous sections, we have discussed two surrogate-modeling-based methods for approximating $\phi$, which follow rigorous function-approximation and computational-linear-algebra theory. In this section, we describe an alternative approach, which approximates $\phi$ following more heuristic arguments. Thanks to some drastic simplifications, the upcoming “hybrid” method, described in Section 5.3 below, is extremely efficient. Indeed, it builds a surrogate model using roughly as few samples as Algorithm 2 (hence, at roughly the same cost), with an extremely low surrogate-evaluation cost $c_{\textup{surrogate}}=\mathcal{O}\left(n_{K}\right)$.
@@ 5.2 Assembling the approximate field amplification
We start by observing that the coefficients $\{\psi_{\lambda}\}_{\lambda\in\widetilde{\Lambda}_{K}}$ can be interpreted as the degrees of freedom appearing in an approximation of $\phi$ that uses a linear combination of the “basis functions”
$\left\{\mathbb{R}\ni k\mapsto\left|k-\lambda\right|^{-1}=\left((k-\operatorname*{Re}\lambda)^{2}+(\operatorname*{Im}\lambda)^{2}\right)^{-1/2}\in\mathbb{R}\right\}_{\lambda\in\widetilde{\Lambda}_{K}}.$
As long as the approximate poles are pairwise distinct and pairwise not complex conjugate, the above basis functions are linearly independent. In this setting, any set of $M_{K}^{\prime}$ distinct collocation points $\{\widetilde{k}_{i}\}_{i=1}^{M_{K}^{\prime}}\subset\mathbb{R}$ is “unisolvent” with this basis, in the sense that the data-fit problem
@@
Remark 5.2.
Since the approximate poles are automatically computed as described in Section 5.1, it may happen that two of them may coincide or be complex conjugate, thus making Eq. 26 ill-posed due to repeated basis functions. This issue is easily solved by pre-processing $\widetilde{\Lambda}_{K}$, removing a copy of each repeated or complex-conjugate approximate pole. Note, in particular, that repeated poles might be a symptom of defects in the problem’s spectrum, cf. Remark 3.5 and ensuing paragraph.
Leveraging common ideas from the radial-basis literature, one may improve both approximation accuracy and numerical stability by placing the collocation points at the locations where the basis functions attain their respective maxima. See, e.g., [34]. In our setting, this happens at $\widetilde{k}_{i}:=\operatorname*{Re}\lambda_{i}$ for $i=1,\ldots,M_{K}^{\prime}$. If any real parts of the approximate poles repeat, one should nudge any repeated collocation points by a small (real) amount.
@@
[…]
@@
By construction, this simple choice yields $\phi(\operatorname*{Re}\lambda)=\widetilde{\phi}_{\max,0}(\operatorname*{Re}\lambda)$ for any $\lambda\in\Lambda_{K}$, as long as the maximum in Eq. 30 is attained at $\lambda$, i.e., whenever no approximate pole “hides” any other approximate pole. In particular, note that the pole pre-processing outlined in Eq. 25 greatly helps in preventing this pole-hiding issue.
@@ 5.3 The overall algorithm
In particular, note that, asymptotically in $N_{h}$, building a hybrid approximation $\widetilde{\phi}\in\{\widetilde{\phi}_{\textup{sum},n_{b}},\widetilde{\phi}_{\max,n_{b}}\}$ is as costly as building a matrix-valued rational approximation via Algorithm 2. However, the evaluation of a hybrid $\widetilde{\phi}$ is $N_{h}$-independent, in contrast to the $\mathcal{O}\left(N_{h}^{2}\right)$-cost of evaluating the surrogate built with Algorithm 2.
This enhanced “online efficiency” (using surrogate modeling jargon) is confirmed in our numerical tests.
@@
[…]
@@ 6.1.1 Spectral features
Moreover, we can see that one of the rational poles, $\lambda\approx 3.16+0.07\textup{i}$, has positive imaginary part. This is a non-physical eigenvalue, but should not be necessarily considered spurious, cf. Remark 5.1. Indeed, rational approximation sometimes relies on unstable poles to improve approximation accuracy. It is commonly suggested to either keep such non-physical poles, to “flip” them [13], i.e., change the sign of their imaginary parts, or to remove them [18]. Four more spurious poles can also be observed in Fig. 2 (top), along a band with imaginary part between 1-1 and 0.2-0.2.
@@ 6.1.2 Field amplification
A partial explanation for this effect can be found by looking at the shape of the $14$ basis functions $\left|\cdot-\lambda\right|^{-1}$, cf. Section 5.2, that are removed when poles are filtered. As a specific example, the basis function corresponding to the spurious pole $\lambda_{1}\approx 1.20-0.62\textup{i}$, namely, $((k-1.20)^{2}+0.62^{2})^{-1/2}$, is more “spread out” than the basis function corresponding to the non-spurious pole $\lambda_{2}\approx 1.22-0.04\textup{i}$, whose real part is similar but whose imaginary part is much smaller, namely, $((k-1.22)^{2}+0.04^{2})^{-1/2}$. Due to its larger variance, the former basis function has a global effect on the approximation $\widetilde{\phi}_{\textup{sum},0}$, breaking the “upper-bound” property of Eq. 24 at wavenumbers far from $\operatorname*{Re}\lambda_{2}$.
We also test the hybrid algorithm with $n_{b}>0$ extra basis elements, by taking a piecewise-linear finite element basis on a uniform partition of $K$ into $(n_{b}-1)$ intervals. In our experiment, we set $n_{o}=n_{b}=M_{K}^{\prime}=20$, the number of (filtered) rational poles. This leads to an enhanced approximation accuracy. Interestingly, the “max” approximation $\widetilde{\phi}_{\textup{max},n_{b}}$, defined in Eq. 30, seems to perform better than the “sum” one $\widetilde{\phi}_{\textup{sum},n_{b}}$, defined in Eq. 27.
@@ 6.2 C-shaped and kite-shaped scatterer
C shape
Algorithm 1 $3.60\cdot 10^{4}$
@@
kite
Algorithm 1 $3.21\cdot 10^{4}$
@@ 7 Conclusions
  • The “hybrid” method from Algorithm 4 targets the field amplification more directly, through a scalar interpolatory approach. This reduces the evaluation cost at the price of some accuracy, while keeping the training cost comparable to that of Algorithm 2. The hybrid method uses rational approximation as a tool to greatly improve the approximation quality and efficiency, by enriching the interpolation basis with some additional rational terms. This strategy ultimately enables a modest approximation accuracy with an evaluation cost that is much lower than that of Algorithm 2.
@@
Concerning future research directions, we note that the field amplification has been defined in terms of surface potentials on the interface $\Gamma=\partial\Omega$. In some cases, it might be of greater interest to consider volume field amplifications, e.g., measured in the $H^{1}(\Omega)$-metric: $\phi_{\Omega}(k):=\sup_{u_{\textup{inc}}\neq 0}\left\|u_{\textup{tot}}(k)\right\|_{H^{1}(\Omega)}/\left\|u_{\textup{inc}}\right\|_{H^{1}(\Omega)}$, cf. Eq. 1. In our setting, $\phi_{\Omega}$ may be computed via the representation formula Eq. 5, which provides an invertible extension operator from surface potentials to Helmholtz solutions over $\mathbb{R}^{d}$. In principle, our discussion and our techniques should apply also to $\phi_{\Omega}$. In particular, our methods should be able to well approximate the locations of the problem’s quasi-resonances, which appear as “peaks” both in $\phi$ and $\phi_{\Omega}$. However, in such case the approximation task is much more challenging, since (i) the wavenumber-dependent extension operator makes $\phi_{\Omega}$ depend on $k$ in a more complicated way than $\phi$, and (ii) building and inverting the extension operator is difficult in itself. For these reason, further analysis and implementation work remain to be done.
@@ Appendix A Exact scattering eigenvalues on the disk
$\textup{find }\lambda\in\mathbb{C}\textup{ and }(\alpha,\beta)\in\mathbb{C}^{2}\setminus\{\mathbf{0}\}\textup{ s.t.}\begin{pmatrix}J_{\nu}(n_{\textup{i}}\lambda)&H_{\nu}(\lambda)\\ n_{\textup{i}}\lambda J_{\nu}^{\prime}(n_{\textup{i}}\lambda)&\lambda H_{\nu}^{\prime}(\lambda)\end{pmatrix}\begin{bmatrix}\alpha\\ -\beta\end{bmatrix}=\mathbf{0}.$